OpenStudy (anonymous):

list all the possible rational and actual rational roots for the equation x^3+2x^2-5x-6=0

sam (.sam.):

First look at the multiples of the constant in x^3+2x^2-5x-6=0, \[[6]=\pm1,\pm2,\pm3,\pm6\] You can try an error by substituting these\[ \pm1,\pm2,\pm3,\pm6\] Into the equation and see whether you get zero. ---------------------------------------------------------------- In this case, x+1 is a factor of x^3+2x^2-5x-6=0, So, (x+1)( Quadratic ) -------------------------------------------------------------- To find the quadratic, use x+1 and do long division, x^2+x-6 _______________________________ x+1) x^3+2x^2-5x-6 x^3+x^2 -------------------------- x^2-5x x^2+x -------------------- -6x-6 -6x-6 --------------- -------------------------------------------------------------- So, (x+1)( Quadratic ) (x+1)( x^2+x-6 ) Factor the quadratic, (x+1)(x-2)(x+3) Roots are x=-1,x=2,x=-3